# Overcoming Data Scarcity in Generative Language Modelling for Low-Resource Languages: A Systematic Review

### v1

@@ 4.7. RQ7: How Scarce Is the Data in Studies Tackling Low-Resource Language Modelling?
The availability of data for building language models differs considerably between LRLs. Furthermore, it is important to discuss that measuring and comparing the scarcity of data presents a challenge given that the literature often reports different metrics relating to the training data. Some papers clearly document the number of sentences used to train models for LRLs (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Hong et al., 2024; Gerz et al., 2018). On the other hand, several studies quantify the data in terms of size (Salim et al., 2023; Downey et al., 2024; Acikgoz et al., 2024; Cahyawijaya et al., 2021). Some papers quantify the size of their training dataset by reporting the number of tokens in their respective datasets (Toraman, 2024; Chang et al., 2023; Maimaiti et al., 2021). Moreover, some studies in this review omit metrics relating to the size of their datasets (Niyogi and Bhattacharya, 2024; Holiel et al., 2023). Overall, these varying methods of reporting dataset metrics highlight a limitation in comparing different language modelling approaches for LRLs, given that there is not a universal metric enlisted by these studies that allows us to evaluate and compare each dataset. Although it is a challenge to assess the scarcity across different studies, some insights can be gained from our analysis. Firstly, we propose that there is a sliding scale of data scarcity for LRLs. Specifically, some LRLs such as Khmer were modelled using three different datasets (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022), whereas many of languages modelled by the studies in this review were only trained on a single dataset (Berckmann and Hiziroglu, 2020; Li et al., 2024b). While some LRLs such as Uyghur were modelled using a variety of datasets by different studies (Mao and Yu, 2024; Mi et al., 2024), other LRLs have been modelled by a single study using much more data (Gerz et al., 2018). These findings suggest that there is a data disparity within the group of LRLs itself. Secondly, we observe that out of the studies that quantify their data in terms of size, the Turkish dataset is significantly larger than the others (Acikgoz et al., 2024). This reflects the conflicting stance on the resource level of Turkish within the literature in this review (Chang et al., 2023; Toraman, 2024; Yirmibeşoğlu and Güngör, 2023). This contributes to our findings that there is a sliding scale of data scarcity, even among the LRLs themselves.
@@ 4.8. RQ8: How Are the Models in These Studies Being Evaluated?
These studies evaluate generative language models for LRLs in many different ways for a broad range of NLG tasks. Similarly to RQ7, the diversity of reported evaluation metrics introduces challenges when comparing results between studies in this review. Given that half of the papers in this review tackle translation, it is unsurprising that BLEU dominates as the most frequently used evaluation metric. However, it is worth highlighting that the BLEU metric suffers from a number of known shortcomings relating to the reporting of results. Post highlights that preprocessing techniques can have a significant impact on BLEU scores, thus impeding on comparability between studies (Post, 2018). Some papers in this review enlist sacreBLEU as a means to address BLEU-based limitations (Wongso et al., 2023; Hong et al., 2024; Liu, 2022). The use of metrics such as ROUGE, chrF, METEOR and relatively newer methods like COMET and sacreBLEU indicates that a wide variety of fine-tuned task-specific models are being trained by studies in this review. The infrequent use of human evaluation methods reveals limitations in most of the studies given that such qualitative insights into model output are important when modelling LRLs. Evaluating whether a model successfully captures the cultural qualities of a language should be essential in determining if an LRL has been modelled accurately. Given that modelling an LRL would be a form of cultural preservation, it would be important to include a qualitative evaluation that ties the research output back to a human context. The lack of a universal approach to evaluating LRL models and the use of flawed and dated, though still prevalent, metrics presents a challenge when comparing the results of different studies within this field of generative language modelling.
@@ 4.9. RQ9: What Is the NLG Task?
While a variety of different NLG tasks are modelled by studies in this review, translation dominates as the most prevalent task. Given that this review focuses on modelling LRLs, this dominance could be attributed to the need for translation models as an initial step in digitising and creating models for these languages. Furthermore, the resources for building translation models could be more readily available in contrast with the extensive resources that are required for training other generative models such as dialogue generation models. This observation could further highlight that LRLs are trailing significantly behind more widely spoken languages such as English in terms of the research output for them. Some of the studies contribute a general language model that can be adapted and fine-tuned for different tasks (Gerz et al., 2018; Salim et al., 2023). This type of contribution could be an initial step that could yield further research relating to downstream generative language tasks. Question-answering was also found to be frequently modelled within this review. This research interest could reflect a demand for information retrieval systems for LRLs. Other NLG tasks such as summarisation and dialogue generation were rarely modelled by studies in this review. This could also highlight that LRLs are lacking in research for these downstream tasks. Given the popularity of generative language models, we would expect the frequency of natural language generation tasks to shift from translation to dialogue generation and general language model research in the future.
@@ 5. Limitations
This section outlines the considerations and limitations relating to this systematic review. Given that LRLs are a central element of this research, excluding non-English papers could omit relevant literature that addresses the challenge of building generative language models with limited data. Furthermore, the title, abstract and keywords of a paper were examined to identify their relevance during the second pass of the study selection phase. This introduces a limitation in this systematic review as relevant information could have been present in the identified papers but absent from their respective metadata. Although both popular platforms and NLP-specific repositories were searched for relevant studies, relevant studies could exist in less conventional, but still significant sources such as niche linguistic journals or conferences. While the inclusion of synonyms for keywords such as ”low resource language” and ”language model” improves coverage, it does not guarantee that all relevant studies will be identified. In terms of performing the study selection, the use of ASReview as a tool to improve the initial screening process could introduce inaccuracies. The tool does not classify studies but instead predicts their relevance using active learning. Therefore, algorithmic bias could be introduced whereby a reviewer could expect studies seen at the beginning of the selection process to be relevant and could consequently include irrelevant studies. It is worth reporting that a single reviewer worked independently on each stage of the screening process. Although this increases the risk of subjectivity and errors, this reflects the challenges involved with sourcing adequate resources such as multiple reviewers to mitigate bias during the study selection phase. The heterogeneity of results across the different studies imposes a limit on the validity of the comparisons and conclusions between them. Many of the studies report the same metrics such as BLEU, however not all of them include a detailed description of their experimental setup. Moreover, some studies report different or updated forms of metrics (sacreBLEU) for the same generative tasks that make it challenging to compare their results accurately.
@@ 6. Conclusion
7. Conflicts of Interest
The authors declare that they have no conflicts of interest related to this work.

### v2

@@ 4.7. RQ7: How Scarce Is the Data in Studies Tackling Low-Resource Language Modelling?
The availability of data for building language models differs considerably between LRLs. Therefore, it is important to highlight that measuring and comparing the scarcity of data presents a challenge given that the literature often reports different metrics relating to the training data. While some papers report the number of sentences used to train models for LRLs (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022; Hong et al., 2024; Gerz et al., 2018), others quantify by size (Salim et al., 2023; Downey et al., 2024; Acikgoz et al., 2024; Cahyawijaya et al., 2021) or by tokens in their respective datasets (Toraman, 2024; Chang et al., 2023; Maimaiti et al., 2021). On the other hand, some studies omit metrics relating to the size of their datasets completely (Niyogi and Bhattacharya, 2024; Holiel et al., 2023). Overall, these varying methods of reporting dataset metrics present a limitation in comparing different language modelling approaches for LRLs, given that there is not a universal metric or framework used by these studies that allows us to evaluate and compare each dataset. Although it is a challenge to assess the scarcity across different studies, our analysis reveals two key insights. Firstly, the findings show that there is a gradient of data scarcity for LRLs. Specifically, some LRLs such as Khmer were modelled using three different datasets (Agarwal et al., 2022; Mao and Yu, 2024; Pham Van and Le Thanh, 2022), whereas many of the languages modelled were only trained on a single dataset (Berckmann and Hiziroglu, 2020; Li et al., 2024b). While some LRLs such as Uyghur were modelled using a variety of datasets by different studies (Mao and Yu, 2024; Mi et al., 2024), other LRLs have been modelled by a single study using much more data (Gerz et al., 2018). These findings suggest that there is a data disparity within the group of LRLs itself.
Secondly, out of the studies that quantify their data in terms of size, the Turkish dataset is significantly larger than the others (Acikgoz et al., 2024). This reflects the conflicting stance on the resource level of Turkish within the literature (Chang et al., 2023; Toraman, 2024; Yirmibeşoğlu and Güngör, 2023) and consequently contributes to the need for a more robust framework for determining the resource level of a given language. Future work should focus on producing a universal framework or series of metrics for reporting on the resource level of a LRL. A universal framework would make comparisons between future LRL-based studies much easier. Additionally, it would enable future researchers to find the most underrepresented LRLs and gauge their suitability for generative modelling based on the available data for a given language. Future work on systematically reporting data resources could contribute to a more comprehensive understanding of cross-lingual resources for modelling multiple languages or language families. Overall, this review finds that the lack of standardisation for reporting available data, combined with the existing data disparity among LRLs, presents a barrier to progress in generative LRL modelling.
@@ 4.8. RQ8: How Are the Models in These Studies Being Evaluated?
These studies evaluate generative language models for LRLs in many different ways due to the broad range of NLG tasks being modelled. Similarly to the challenges of quantifying dataset size in RQ7, the diversity of reported evaluation metrics introduces challenges when comparing results between studies. Given that half of the papers in this review tackle translation, it is unsurprising that BLEU dominates as the most frequently used evaluation metric. However, it is worth highlighting that the BLEU metric suffers from several known shortcomings relating to the reporting of results. Post highlights that preprocessing techniques can have a significant impact on BLEU scores, thus impeding comparability between studies (Post, 2018). The use of sacreBLEU as a means to address BLEU-based limitations (Wongso et al., 2023; Hong et al., 2024; Liu, 2022) and metrics such as ROUGE, chrF, METEOR and relatively newer methods like COMET and sacreBLEU indicates that a wide variety of fine-tuned task-specific models are being trained.
However, the infrequent use of human evaluation methods reveals considerable limitations in most of the studies. Given that modelling an LRL contributes to cultural preservation, evaluating whether a model successfully captures the qualities of a language should be essential in determining if an LRL has been modelled accurately. Therefore, in the interest of ethical language modelling, studies should not prematurely report success when modelling a LRL without a degree of human evaluation, as the statistical metrics do not systematically quantify whether a model for a LRL is representative of it or not. The lack of a universal approach to evaluating LRL models and the use of flawed and dated, though still prevalent, metrics presents a challenge when comparing the results of different studies within this field of generative language modelling. Future research should focus on building a framework for evaluating LRL models across the spectrum of generative tasks. Additionally, future work should endeavour to include human, qualitative evaluation when evaluating a model’s capacity to generate text in a given LRL.
@@ 4.9. RQ9: What Is the NLG Task?
Given that this review focuses on modelling LRLs, the prevalence of translation may be attributed to the need for translation models as an initial step in digitising and creating other generative models for these languages. Furthermore, the resources for building translation models are likely more readily available in contrast with the extensive resources that are required for training other generative models such as dialogue generation models. The predominance of translation suggests that LRLs lag significantly behind widely spoken languages such as English, both in research attention and in the availability of supporting resources. On the other hand, studies contributing a general language model that can be adapted and fine-tuned for different tasks (Gerz et al., 2018; Salim et al., 2023) could be an initial step that would yield further research relating to downstream generative language tasks.
Additionally, the prevalence of question-answering models likely reflects a demand for information retrieval systems for LRLs. Given that other NLG tasks such as summarisation and dialogue generation are rarely modelled, it is clear that LRLs are lacking in research for downstream beyond translation and QA tasks. This is likely due to the lack of large, complex and task-specific labelled corpora required to train such downstream models. Furthermore, a pre-trained model is typically required as a base for fine-tuning task-specific models. Therefore, future research should be focused on releasing open-source pre-trained models for LRLs, as this would enable researchers to tackle other downstream generative tasks. Given the current popularity of generative language models, it is likely that the frequency of NLG tasks will shift from translation to general language model research and downstream generative tasks in the future.
However, it is worth considering that, in theory, a strong translation model could be the key to unlocking generative technology access for LRL communities. More specifically, an effective translation model could interface with the input and output of a better-resourced, higher-performing model for majority languages, such as English. This could be a sustainable way of reducing language inequality in the digital world as such translation models could wrap around future state-of-the-art models with relative ease. Future work should investigate the feasibility and quality of output when wrapping existing generative models with a LRL translation model.
@@ 5. Limitations
This section outlines the considerations and limitations relating to this systematic review. Given that LRLs are a central element of this research, excluding non-English papers could omit relevant literature that addresses the challenge of building generative language models with limited data. Furthermore, the title, abstract and keywords of a paper were examined to identify their relevance during the second pass of the study selection phase. This introduces a limitation in this systematic review as relevant information could have been present in the identified papers but absent from their respective metadata. Although both popular platforms and NLP-specific repositories were searched for relevant studies, relevant studies could exist in less conventional, but still significant sources such as niche linguistic journals or conferences. While the inclusion of synonyms for keywords such as ”low resource language” and ”language model” improves coverage, it does not guarantee that all relevant studies will be identified. In terms of performing the study selection, the use of ASReview as a tool to improve the initial screening process could introduce inaccuracies. The tool does not classify studies but instead predicts their relevance using active learning. Therefore, algorithmic bias could be introduced whereby a reviewer could expect studies seen at the beginning of the selection process to be relevant and could consequently include irrelevant studies. It is worth reporting that a single reviewer worked independently on each stage of the screening process. Although this increases the risk of subjectivity and errors, this reflects the challenges involved with sourcing adequate resources such as multiple reviewers to mitigate bias during the study selection phase. The heterogeneity of results across the different studies imposes a limit on the validity of the comparisons and conclusions between them. Many of the studies report the same metrics such as BLEU, however not all of them include a detailed description of their experimental setup. Moreover, some studies report different or updated forms of metrics (sacreBLEU) for the same generative tasks, making accurate comparisons of their results challenging.
@@ 6. Conclusion
7. Statements and Declarations
Conflict of Interest The authors declare that they have no conflicts of interest related to this work.